MMCC/HtX Reference Card

(last update: June 8, 1997)

1 Document Structure

\documentstyle{mmcc}
\me{...}
\lib[ffffff]{vcbl}{Vocabulary}
\wallpaint{....}
\begin{document}
\docheader{<Course Name>}{<Module #>: <Module name>}
\begin{goallist}
\item ..
\item ..
\end{goallist}
\begin{prereq}
\item ..
\item ..
\end{prereq}
<Introductory material>
\section{..}
\subsection{...}
\vocab   % if necessary
\doindex % if necessary
\refstart % if necessary
\refentry{..} ...
\refentry{..} ...
\refend
\end{document}

2 Visual Style

2.1 Typeface

italics  {\it ... }
boldface  {\bf ...}

typewriter {\tt ...}

2.2 Lists

ordered list  \begin{enumerate} \item .. \item .. \end{enumerate}
unordered list  \begin{itemize} \item .. \item .. \end{itemize}
description/definition \begin{definition} \item[..] .. \item[..] .. \end{enumerate}

2.3 Tables

unlined \begin{tabular}{...} ... & ... & ... \\ ... & ... \\ ... \end{tabular}
lined  \begin{tabular}{|..|..|} ... & ... & ... \\ ... & ... \\ ... \end{tabular}

2.4 Font Size

absolute {\font{n} ...} (n = 0 … 5)
relative {\font{±n} ...} (n = 0 … 5)

2.5 Headers

HTML header \header{n} (n = 1 … 5)

2.6 Display

paragraph  \par (or empty line)

centering  \begin{center} ... \end{center}

                    {\centering ... }

quotation  \begin{quote} ... \end{quote}
HTML pre-formatted \begin{verbatim} ... \end{verbatim}
code  \begin{code} ... \end{code}

                    {\code ... }

2.7 Math Mode

in-text formula  $ ... $
displayed formula $$ ... $$
subscript  _{...}
superscript  ^{...}

fraction  \frac{..}{..}

ellipsis  \ldots

3 HTML features

3.1 Images

simple image  \image{<filename>}
image with alignment \image[<left|center|right>]{<filename>}
image with alt title  \image[<left|center|right>][<alt title>]{<filename>}

3.2 Anchors and Links

external link \xlink{<URL>}{<hot item>}
anchor  \anchor{<label>}{<hot item>}
internal link \link{<label>}{<hot item>}

3.3 Extensibility

HTML command  \mkp*[<options>]{tag} (becomes <tag options>)
HTML environment {\mkp[<options>]{tag} ... } (becomes <tag options> ... </tag>)

3.4 Footnotes/Glossary

\begin{gloss*}[<title of entry>]{<hot item>} <entry text> \end{gloss*}

3.5 Reference Citations

citations in text \cite{<keyword>}{<hot item>}
references  \refstart \refentry{..} ... \refentry{..} ... \refend

               \refentry{<keyword>} <entry>

3.6 Indexing

index annotation using standard icon  \indx{<index keyword>}
index annotation using specified hot text \indx[<hot text>]{<index keyword>}
index creation  \doindex

3.7 Vocabulary

vocabulary annotation in text \vref{<word>}

                           \vref[<keyword>]{<word>}

Requires auxiliary vocabulary file.

3.8 Pedagogical Features

questions & answers \begin{quest} ... \ans ... \end{quest}
multiple choice  \begin{mcquest} ... \mans ... \mans* ... \end{mcquest}
exercise  \begin{exercise} ... \end{exercise}
exercise solution  \begin{exsol} ... \end{exsol}
hint  \begin{hint}{<hot item>} .. \end{hint}

3.9 HTML Display

\begin{html-display} ... \end{html-display}